OpenStudy (anonymous):

Simplify the expression. 55 – 4 • 5 + 12

OpenStudy (anonymous):

No answer chooses for this problem

OpenStudy (kendricklamar2014):

\[4*5 = ?\] @THEHELPER123

OpenStudy (anonymous):

20

OpenStudy (kendricklamar2014):

\[55-20=?\]

OpenStudy (anonymous):

35

OpenStudy (kendricklamar2014):

\[35+12=?\]

OpenStudy (anonymous):

47

OpenStudy (kendricklamar2014):

\[Answer = 47\]

OpenStudy (anonymous):

thank you and i gave u a medal

OpenStudy (kendricklamar2014):

No Problem and Thank You